OpenStudy (anonymous):

find the slope and y intercept 7-1/5y=5x that is 7 minus 1 over 5 y =5x

OpenStudy (chaise):

This is a hyperbola - the graph is constantly changing slopes.

OpenStudy (anonymous):

it is ask for the slop and y intercept

OpenStudy (anonymous):

This is easiest to do if you put it into slope-intercept form, y=mx+b. 7 - (1/5)y = 5x -(1/5)y = 5x - 7 (1/5)y = -5x + 7 y = -25x + 35 m is your slope, and b is the intercept.